OpenStudy (anonymous):

When dividing the complex numbers, the first step is to multiply top and bottom by the complex _____ f the denominator. Is the correct answer conjugate?

OpenStudy (anonymous):

Yep :)

OpenStudy (anonymous):

Thank you so much!:)

OpenStudy (anonymous):

No problem. Easiest "help" I've ever given!

OpenStudy (anonymous):

Could you help me with one more thing?:)

OpenStudy (anonymous):

\[\frac{ 2+3i }{ 1+2i }\]

OpenStudy (anonymous):

Ok. What's the conjugate of the denominator?

OpenStudy (anonymous):

2??

OpenStudy (anonymous):

When we take a complex conjugate, we just change the sign in front of everything with an "i" \[(a+b+c+di)^* = (a+b+c-di)\]

OpenStudy (anonymous):

oh yes! Okay I remember now.. Got it.

OpenStudy (anonymous):

So from there what would I do?

OpenStudy (anonymous):

Multiply both the top and the bottom by the conjugate of the denominator

OpenStudy (anonymous):

It should look like: \[\frac{2+3i}{1+2i}\frac{1-2i}{1-2i} = ?\]

OpenStudy (anonymous):

\[\frac{ -4-i }{ 3}\]

OpenStudy (anonymous):

Isn't that the final answer?

OpenStudy (anonymous):

That's not what I get. Let me double check mine

OpenStudy (anonymous):

wait! no it's over 5!

OpenStudy (anonymous):

Closer to what I get :)

OpenStudy (anonymous):

(2+3i)(1-2i) = 2*1+3i*1+2*(-2i) + 3i*(-2i)

OpenStudy (anonymous):

Oh, I see what happened on yours. 3i * -2i = -6*i^2 = 6

OpenStudy (anonymous):

The denominator is 5 isn't it?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

Okay got it.. Thought so.

OpenStudy (anonymous):

What is your final answer? Your numerator wasn't quite correct, before.

OpenStudy (anonymous):

Wait wait wait! is the top 8-i!?

OpenStudy (anonymous):

There you go :)

OpenStudy (anonymous):

\[\frac{ 8-i }{ 5 }\]

OpenStudy (anonymous):

That's what I get :) Good work!

OpenStudy (anonymous):

Thank you so much for all the help!

OpenStudy (anonymous):

My pleasure. Keep up the good work!